A company issues term bonds totaling $300,000 on January 1, 2014. The bonds have a coupon rate of 5%, pay interest semi-annually on January 1st and July 1st of each year, and mature in 10 years.

Calculate the bond issue price assuming that the prevailing annual market rate of interest is: 5%, 4%

As applicable, prepare a bond discount or bond premium amortization schedule based on the effective-interest method.

As applicable, record the applicable journal entries in 2014 (1/1/2014, 7/1/14, 12/31/14).

1) 5% Annual market rate of interest (at par) Since the coupon rate equals the prevailing market rate (both 5%), there is no bond discount or premium. And since there is no bond discount or premium, there is no need for an amortization table.

N= I/Y= PV= PMT= FV= The bond issue price equals:

Semi-annual interest is equal to:

Applicable journal entries (first year, 1/1/2014, 7/1/14, 12/31/14):

2) 4% Annual market rate of interest (Bond premium) In this scenario, the coupon rate of 5% is greater than the prevailing market rate of 4%. Therefore, this bond will be issued at premium.

N= I/Y= PV= PMT= FV= The bond issue price equals:

Amortization Schedule:

3) 6% Annual Market rate of interest (bond discount) In this scenario, the coupon rate of 5% is less than the prevailing market rate of 6%. Therefore, this bond will be issued at a discount.

N= I/Y= PV= PMT= FV= The bond issue price equals:

Amortization Schedule:
Answer #1

1. Ifannual market rate of interest is 5%, then it means it is equal to coupon rate which means bond will be issued ar par i.e. bond will be issued for $300000 and this there will be no requirement for bond amortization table or schedule and journal entries in this case will be as follows :-

Date general journal debit credit
1/1/2014

cash

Bonds payable

(To bond issued at par value )

300000

300000

7/1/2014

interest exp

Cash

(Interest on bond payable semi annually as 300000x5%x6/12)

7500

7500

12/31/14

interest exp

Cash

(To interest payable on bonds semi annually )

7500

7500

2. Since market rate of interest is 4% and coupon rate of bond is 5% , hence bonds will be issued at premium. Bond price will be calculated as follows

7500xpvaf(2%, 20)+300000xpvif (2%,20)

7500x16.3514+300000x0.6730

= 122635+201900

= $324535.

Amortization schedule will be as follows :-

(A)

Period

(B)

interest payment (coupon)

(C= market rate previous bv in g)

interest expense (market rate)

(D = c-b)

amortization of bond premium

(E)

balance in premium

(F)

balance in bond payable

(G = F+E)

book value of bonds

1/1/14 24535 300000 324535
7/1/14 7500 6491 (1009) 23526 300000 323526
12/31/14 7500 6471 (1029) 22497 300000 322497

Journal entries will be as follows :-

Date general journal debit credit
1/1/14

cash

Bonds payable

Premium on bonds payable

(To bonds issued at premium)

324535

300000

24535

7/1/14

interest expense


Premium on bonds payable

Cash

(To bonds interest paid)

6491

1009

7500

12/31/14

interest exp

Premium on bonds payable

Cash

(To interest on bonds paid)

6471

1029

7500

3 . If market rate of interest is 6%, it means bonds has to be issued at discount. Bond price will be calculated as follows :-

7500xpvaf (3%,20)+300000xpvif (3%, 20)

= 7500x14.8775+300000x0.5537

= 111581+166110

= $277690.

Discount on issue of bond = 300000-277690 = $22310

Amortization schedule will be as follows :-

(A) to

Period

(B = 2.5% of fv)

interest payment

(C = 3% of previous by)

Interest expense

(D = C-B)

Amortization of bond discount

(E)

Balance in bond discount

(F)

Balance in bond payable

(G = F-E)

Book value of bond

1/1/14 22310 300000 277690
7/1/14 7500 8330 830 21480 300000 278520
12/31/14 7500 8355 855 20625 300000 279375

Journal entries will be as follows :-

Date general journal debit credit
1/1/14

Cash

Discount on bonds payable

Bonds payable

(To bonds issued at discount(

277690

22310

300000

7/1/14

interest exp

Discount on bonds payable

Cash

(To interest on bonds paid)

8330

830

7500

12/31/14

interest exp

Discount in bonds payable

Cash

(To interest on bonds paid)

8355

855

7500

Note :-

Kindly note that there may be minute rounding off problem. Please take care of it.